# GPU-Accelerated Parallel Selected Inversion for Structured Matrices Using sTiles

### v1

@@ I Introduction
To address this challenge, selected inversion techniques have emerged as a focused alternative, computing only specific entries or substructures of the inverse matrix that are required for the target application. By avoiding the computation of elements considered unnecessary, the selected inversion significantly reduces resource consumption. This makes it particularly valuable in domains such as electronic structure calculations [1], large-scale sparse inverse covariance estimation [2], and Bayesian modeling [3]. However, despite these advantages, achieving scalability in selected inversion methods remains a significant challenge, particularly for high-dimensional applications.
The sparsity patterns of arrowhead matrices offer unique computational advantages, particularly in Bayesian inference, where they model interactions between multiple random effects. These matrices are commonly constructed using Kronecker products, for example, by capturing space-time dependencies in spatio-temporal models. Figure 1 illustrates this by comparing two arrowhead matrices (matrix A and matrix B) with differing sparsity levels. The sparsity level of these matrices often reflects the number of spatial locations or time points used in a model. As data collection technologies advance, we are entering an era of increasingly high-resolution datasets, where more sensors, satellites, and monitoring systems generate vast amounts of spatial and temporal information, leading to lower levels of sparsity in arrowhead matrices, driven by higher resolution and wider matrix bandwidth, as the number of locations increases.
@@
To our knowledge, there is limited literature on parallel selected inversion methods specifically designed for sparse symmetric matrices with an arrowhead structure that utilize tile-based computations and the $LL^{T}$ Cholesky factorization. Existing methods, such as the Supernodal Selected Inversion (SINV) algorithm [9], compute selected inverse elements by leveraging supernodal partitions from LU factorizations and traversing elimination trees in parallel. Similarly, the SelInv algorithm [10] targets sparse symmetric matrices and employs $LDL^{T}$ factorizations with block algorithms and supernodes for efficiency. While both approaches benefit from multithreaded BLAS-level parallelism, they face challenges with irregular memory access patterns and dependency management inherent to supernodes.
Additional advancements include the selected inversion method for block tridiagonal arrowhead (BTA) matrices introduced in [11], which builds on strategies from quantum transport simulations and Kalman-Bucy filtering to efficiently compute block diagonal elements. However, its reliance on the specific BTA structure limits its general applicability and parallelization potential. Likewise, the hybrid parallel approach in [3] combines Krylov subspace methods and domain decomposition to approximate selected inverse elements in spatio-temporal Gaussian Markov Random Fields (GMRFs) [12]. Although scalable, this method relies on approximations through sampling and heuristic parameters.
Parallel implementations of Takahashi-based selected inversion for general sparse matrices have been explored, including a distributed-memory approach by [13], which integrates block-based Schur complement decomposition and relies on sparse direct solvers (e.g., PARDISO [18]) for symbolic and numeric factorization. In contrast, our method targets structured sparse symmetric matrices (e.g., arrowhead form), and achieves exact parallel solutions through tile-based computations, static scheduling, and GPU acceleration.
@@
Building on this foundation, tile methods provide finer granularity, enhanced data locality, and superior scalability, making them especially effective for structured matrices like arrowheads. sTiles [14], a tile-based framework for high-performance linear algebra, exemplifies these advantages by utilizing sparse-dense tile computations for efficient factorizations and subsequent matrix operations. By design, sTiles minimizes inter-task dependencies and optimizes parallel execution, forming a robust foundation for extending selected inversion techniques. In this work, we extend the functionality of sTiles to efficiently perform selected inversion, significantly broadening its practical applicability to broader matrix structures.
@@
This article is organized as follows. Section 2 introduces a recursive tile-based inversion algorithm built upon Cholesky decomposition. In Section 3, we detail a two-phase parallel algorithm, highlighting its static scheduling approach and GPU implementation with a focus on arrowhead structures. Section 4 presents performance results, including comparisons with existing libraries and the effectiveness of GPU implementation. Finally, we conclude with a summary of our findings and discuss potential future directions.
II Recursive Tile-Based Inversion via Cholesky Decomposition
@@
Efficient inversion of large symmetric positive-definite matrices is central to many scientific applications, particularly when only selected entries of the inverse are needed. By leveraging the Cholesky factorization and operating at the tile level, we enable an approach that exploits data locality, facilitates parallelism, and scales well across hardware platforms.
@@
where $L$ is a lower triangular matrix, forms the basis for many matrix inversion algorithms. Traditional approaches compute the inverse of $A$ by inverting $L$ elementwise and then computing $(L^{-1})^{T}$, but these methods are often inefficient for large matrices. To overcome these limitations, tile-based algorithms decompose $L$ into smaller blocks or tiles, enabling parallel computation of the inverse while managing interdependencies between tiles to maintain accuracy.
@@
$L^{T}=\begin{bmatrix}L_{00}&L_{10}&L_{20}\\ 0&L_{11}&L_{21}\\ 0&0&L_{22}\end{bmatrix},\quad\Sigma=\begin{bmatrix}\Sigma_{00}&\Sigma_{01}&\Sigma_{02}\\ \Sigma_{10}&\Sigma_{11}&\Sigma_{12}\\ \Sigma_{20}&\Sigma_{21}&\Sigma_{22}\end{bmatrix},$
@@
Next, we compute $\Sigma_{12}$ using:
$\Sigma_{21}L_{11}+\Sigma_{22}L_{21}^{T}=0,\quad\Sigma_{21}=-\Sigma_{22}L_{21}^{T}L_{11}^{-1}.$
@@
$\begin{array}[]{rl}\Sigma_{22}&=L_{22}^{-T}L_{22}^{-1},\\[6.0pt] \Sigma_{21}&=-\Sigma_{22}L_{21}^{T}L_{11}^{-1},\\[6.0pt] \Sigma_{11}&=-\Sigma_{12}L_{21}^{T}L_{11}^{-1}+L_{11}^{-T}L_{11}^{-1},\\[6.0pt] \Sigma_{20}&=-\Sigma_{21}L_{10}^{T}L_{00}^{-1}-\Sigma_{22}L_{20}^{T}L_{00}^{-1},\\[6.0pt] \Sigma_{10}&=-\Sigma_{11}L_{10}^{T}L_{00}^{-1}-\Sigma_{12}L_{20}^{T}L_{00}^{-1},\\[6.0pt] \Sigma_{00}&=-\Sigma_{01}L_{10}^{T}L_{00}^{-1}-\Sigma_{02}L_{20}^{T}L_{00}^{-1}+L_{00}^{-T}L_{00}^{-1}.\end{array}$
@@
Initialization:
int i, j, k;
for $i=N-1$ to 00 step $-1$ do
     for $j=N-1$ to $i$ step $-1$ do
         if $i==j$ then
              $\Sigma_{ii}\leftarrow\Sigma_{ii}+L_{ii}^{-T}L_{ii}^{-1}$
              for $k=i+1$ to $N$ do
                  $\Sigma_{ii}\leftarrow\Sigma_{ii}-\Sigma_{ik}L_{ki}L_{ii}^{-1}$
              end for
         else
              $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jj}L_{ji}L_{ii}^{-1}$
              for $k=i+1$ to $N$ do
                  if $j!=k$ then
                       $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jk}L_{ki}L_{ii}^{-1}$
                  end if
              end for
         end if
     end for
end for
The core tile operations required for recursive inversion are implemented using well-defined linear algebra kernels. These operations are as follows:
@@
  • LAUUM (Lower Triangular Matrix Multiplication): This kernel updates the diagonal tile by computing the product of a lower triangular matrix and its transpose. The result is stored in the upper triangular part and then mirrored to the full tile:
    $\Sigma_{ii}\leftarrow\Sigma_{ii}\Sigma_{ii}^{T}.$
@@
This targeted computation is particularly beneficial for structured matrices, such as arrowhead or banded forms, where sparsity patterns can be exploited to further reduce overhead. Our approach adopts a static scheduling strategy and targets hybrid CPU/GPU systems for optimal performance. We also generalize the tile-based inversion framework introduced by [11] to support broader classes of structured matrices.
III Parallel Selected Inversion
This section introduces an algorithm for inverting symmetric matrices, with a particular emphasis on arrowhead structures, using a tile-based approach grounded in the Cholesky decomposition $A=LL^{T}$. This Cholesky-based methodology consists of three main phases: heuristic ordering to optimize the preservation of sparsity, symbolic factorization to determine the structure and dependencies within the factorization, and numerical factorization to compute the actual values of the matrix elements. Further details on the Cholesky-based implementation are provided in the context of sTiles [14].
Figure 2: Illustration of Cholesky factorization and inversion patterns of selected tiles for dense and arrowhead matrices
Building on this Cholesky-based framework, the matrix inversion process is structured as follows:
@@ III Parallel Selected Inversion
    The selection step begins with the user specifying a list of matrix elements, identified by their indices $(i,j)$, for which the inverse is required. These indices are mapped to their corresponding tiles using the same compressed tile format employed during the Cholesky decomposition. This step ensures consistency in data representation and prepares the structure for subsequent computations.
@@ III-A Selected Elements of the Inverse
In cases 1-3, the original matrix is fully dense, and the selected tiles include the diagonal. As a result, the inversion covers the entire lower triangle, producing a full inverse. This behavior reflects the high computational cost of inverting dense matrices when the diagonal is involved. Case 6 mirrors this behavior in the arrowhead setting. Although the structure is sparse, selecting the entire matrix leads to full inversion, analogous to case 1.
@@
III-B Directed Acyclic Graph
Figure 3: Directed acyclic graph (DAG) representations for selected inversion on matrices of size $6\times 6$ tiles. The upper plot corresponds to Case 2 (dense matrix) in Figure 2, while the lower plot corresponds to Case 7 (arrowhead matrix). These DAGs visualize task dependencies and parallelism within each structure.
@@ III-B Directed Acyclic Graph
The Directed Acyclic Graph (DAG) captures dependencies between computations, enabling efficient parallel execution and optimized resource allocation. By organizing tasks with well-defined precedence, the DAG ensures that independent computations can proceed concurrently. The DAG underlying our approach is constructed using four key computational kernels—TRSM, LAUUM, GEMM, and TRMM—each applied to specific tiles $\Sigma_{i,j}$, corresponding to the tile at position $(i,j)$. The definitions and roles of these operations are described in detail in Section II.
@@
As illustrated in Figure 3, the DAGs for full and arrowhead matrix inversions (Cases 2 and 7, respectively) exhibit notable differences in structure and parallelism. The DAG for full matrix inversion is characterized by a large width, indicating many tasks that could, in principle, be executed concurrently. In the case of the arrowhead structure, the number of nodes is significantly reduced due to the pruning of unnecessary computations when only selected tiles are targeted for inversion. This leads to a more compact DAG with a level of concurrency depending on which selected tile inversions are needed and reduced overall computational workload. Importantly, the critical path length remains the same in both cases, consisting of six sequential operations along the longest dependency chain for the studied structured matrices.
III-C Two-Phase Algorithm for Selected Inversion
@@
To parallelize Algorithm 1, we adopt a two-phase approach designed to minimize interdependencies between cores, thereby reducing idle times and enhancing parallel efficiency. Beyond parallelism, this approach also prunes unnecessary computations by identifying and skipping tiles that do not contribute to the user-specified subset of the inverse. By leveraging the structure of the selected tiles, we avoid redundant operations while preserving correctness. The division of the algorithm into distinct phases enables better task organization and facilitates scalable execution. A critical aspect of this strategy is the use of static load balancing, where tasks are preassigned to cores during preprocessing. This ensures an even distribution of the workload and significantly reduces runtime scheduling overhead.
Phase 1: Independent Tile Computations
In this phase, each core operates independently on specific columns of the matrix, determined by its thread ID and the total number of participating cores. The columns are processed in a round-robin manner, with the column index assigned based on the core’s thread ID modulo the total number of cores. This approach ensures a balanced and evenly distributed workload across all cores. During this phase, the values in the upper triangular matrix $L^{T}$ are updated to prepare for the inversion in the second phase. Additionally, the intermediate results of the selected inversion are progressively stored in the matrix $\Sigma$, which will eventually contain the selected inverse.
@@
To define the structure, we introduce the notation $j\in\text{neighbors}(i)$ if and only if tile $X_{ij}^{T}$ or tile $X_{ji}^{T}$ ($X=L$ or $X=\Sigma$ ) corresponds to a non-zero tile. A non-zero tile may initially be zero but is added during the symbolic inversion phase and is fully updated by the end of the computation. The algorithm for this phase is detailed in Algorithm 2.
@@
Given: Matrix $A$ is partitioned into $N\times N$ tiles and expressed as $A=LL^{T}$, where $L^{T}$ is the upper triangular factor.
Note: $\Sigma$ is the matrix where the selected inverse is stored.
Initialization: $i\leftarrow N-1-\texttt{thread ID}$
while $i\geq 0$ do
     for all $j\in\text{neighbors}(i)$ and $i\leq j<N-1$ do
         if $i==j$ then
              $\Sigma_{ii}^{T}\leftarrow\texttt{TRSM}(L_{ii}^{T},I)$
         else
              $L_{ij}^{T}\leftarrow\texttt{TRMM}(L_{ii}^{T},L_{ij}^{T})$
         end if
     end for-
[…]
end while
@@
Once all cores have completed their assigned tasks in phase 1, the algorithm transitions to phase 2, where the intermediate results in $\Sigma$ are further processed to compute the final selected inverse.
In phase 2, the intermediate matrix $\Sigma$ is updated using a parallel, task-based approach that respects data dependencies across cores. Each core operates on a specific column $i$, determined by its thread ID, and iterates through dependent rows $j$ and neighboring blocks $k$. Diagonal blocks are updated using contributions from their neighbors, while off-diagonal blocks are computed based on interdependencies between neighboring tiles. The execution follows an asynchronous model, where each core advances independently based on the readiness of required data. A progress tracking mechanism, core_progress, provides lightweight, fine-grained synchronization to enforce correct task ordering without requiring global barriers. This design enables efficient parallelization while handling the complex data dependencies inherent to the selected inversion process.
@@
Initialization: $i\leftarrow N-1-\texttt{thread ID}\texttt{ and }\texttt{set = false;}$;
while $i\geq 0$ do
     for $j=N-1$ to $i$ step $-1$ do
         if $i==j$ then
              $\Sigma_{ii}\leftarrow\texttt{LAUUM}(\Sigma_{ii})$
              for all $k\in\text{neighbors}(i)$ and $i<k<N-1$ do
                  $\Sigma_{ii}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{ik}^{T},\Sigma_{ii})$
              end for
              Set core_progress[i,i] = 1
         else
              if $(i+1)\leq(N-1)$ then
                  set = true; ii = 0; jj = 0;
              end if
              for all $k\in\text{neighbors}(j)$ and $i<k<N-1$ do
                  if $i\in\text{neighbors}(j)\cap j\in\text{neighbors}(k)$ and $i<j$ then
                       if $k>j$ then
                           while core_progress[j,k] $\neq$ 1 do
                                /* Wait */
                           end while
                           $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{kj},\Sigma_{ij})$
                           ii = i; jj = j;
                       else
                           while core_progress[k,j] $\neq$ 1 do
                                /* Wait */
                           end while
                           $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{jk}^{T},\Sigma_{ij})$
                           ii = i; jj = j;
                       end if
                  end if
              end for
              if set then
                  Set core_progress[ii,jj] = 2;
              end if
         end if
     end for
[…]
end while
Figure 4: Directed acyclic graphs for full and arrowhead matrix inversions using 2 cores and 4 cores, with each color representing tasks for each core.

### v2

@@ I Introduction
To address this challenge, selected inversion techniques have emerged as a focused alternative, computing only specific entries or substructures of the inverse matrix that are required for the target application. By avoiding the computation of elements considered unnecessary, the selected inversion significantly reduces resource consumption. This makes it particularly valuable in domains such as large-scale sparse inverse covariance estimation [1], electronic structure calculations [2], and Bayesian modeling [3]. However, despite these advantages, achieving scalability for high-dimensional problems in selected inversion methods remains a significant challenge.
The sparsity patterns of arrowhead matrices offer unique computational advantages, particularly in Bayesian inference, where they model interactions between multiple random effects. These matrices are commonly constructed using Kronecker products, for example, by capturing space-time dependencies in spatio-temporal models [4]. Figure 1 illustrates this by comparing two arrowhead matrices (matrix A and matrix B) with differing sparsity levels. The sparsity level of these matrices often reflects the number of spatial locations or time points used in a model. As data collection technologies advance, we are entering an era of increasingly high-resolution datasets, where more sensors, satellites, and monitoring systems generate vast amounts of spatial and temporal information, leading to lower levels of sparsity in arrowhead matrices, driven by higher resolution and wider matrix bandwidth, as the number of locations increases.
@@
To address the challenges of scalability and hardware efficiency, we build upon the tile-based paradigm. Tile-based methods provide finer granularity, enhance data locality, and expose parallelism for superior scalability, making them especially effective for structured matrices like arrowheads. sTiles [13], a tile-based framework for high-performance linear algebra, exemplifies these advantages by utilizing sparse-dense tile computations for efficient factorizations and subsequent matrix operations. By design, sTiles minimizes inter-task dependencies and optimizes parallel execution, forming a robust foundation for extending selected inversion techniques. In this work, we extend the functionality of sTiles to efficiently perform selected inversion, significantly broadening its practical applicability to broader matrix structures.
@@
The remainder of this article is organized as follows. Section II reviews related work in selected inversion. Section III introduces our parallel tile-based algorithm, beginning with its conceptual foundation in recursive inversion and detailing the two-phase implementation for both CPUs and GPUs. Section IV presents a comprehensive performance evaluation, including comparisons with state-of-the-art libraries and an analysis of GPU acceleration. Finally, Section V concludes with a summary of our findings and discusses potential future directions.
II Related Work
Our work uses a direct, factorization-based method for selected inversion, building upon the foundation laid by Takahashi’s formula [14]. Modern high-performance solvers in this domain, such as MUMPS [15] and PSelInv [16], typically employ supernodal or multifrontal techniques. These methods group columns with similar sparsity patterns into “supernodes” to leverage high-performance BLAS-3 kernels, but their irregular data structures can be challenging to optimize for massively parallel hardware.
Among state-of-the-art solvers for shared-memory systems, Panua-PARDISO [17] is a highly optimized library that uses a robust multi-level parallelization scheme. A comprehensive analysis by Verbosio [18] demonstrated that Panua-PARDISO is significantly faster and more memory-efficient than competitors like PSelInv for the class of sparse-dense problems relevant to our work. These findings establish Panua-PARDISO as the state-of-the-art and our primary benchmark.
In contrast to the supernodal approach, our algorithm is tile-based. We impose a uniform grid of fixed-size blocks (tiles) onto the matrix, creating a highly regular data structure and a predictable dependency pattern. This regularity allows us to employ a static scheduling strategy that maximizes data locality and minimizes runtime overhead, making our approach exceptionally well-suited for structured matrices and hybrid CPU-GPU architectures.
III Parallel Selected Inversion
Our selected inversion algorithm is built upon the Cholesky factorization of a symmetric positive-definite matrix, $A=LL^{T}$. The inverse $\Sigma=A^{-1}$ is computed by solving the block-wise matrix equation $L^{T}\Sigma=L^{-1}$ recursively. This tile-based approach decomposes the problem into a sequence of high-performance operations (e.g., TRSM, LAUUM, GEMM, and TRMM) on small, dense tiles, which enhances data locality and exposes fine-grained parallelism.
III-A Recursive Tile-Based Inversion via Cholesky Decomposition
@@
Efficient inversion of large symmetric positive-definite matrices is central to many scientific applications, particularly when only selected entries of the inverse are needed. By leveraging the Cholesky factorization and operating at the tile granularity, we enable an approach that exploits data locality, promotes parallelism, and scales well across hardware platforms.
@@
where $L$ is a lower triangular matrix, forms the basis for many matrix inversion algorithms. Traditional approaches compute the inverse of $A$ by inverting $L$ element-wise and then computing $(L^{-1})^{T}$, but these methods are often inefficient for large matrices. To overcome these limitations, tile-based algorithms decompose $L$ into smaller blocks or tiles, enabling parallel computation of the inverse while managing interdependencies between tiles to maintain accuracy.
@@
$L^{T}=\begin{bmatrix}L_{00}^{T}&L_{10}^{T}&L_{20}^{T}\\ 0&L_{11}^{T}&L_{21}^{T}\\ 0&0&L_{22}^{T}\end{bmatrix},\quad\Sigma=\begin{bmatrix}\Sigma_{00}&\Sigma_{01}&\Sigma_{02}\\ \Sigma_{10}&\Sigma_{11}&\Sigma_{12}\\ \Sigma_{20}&\Sigma_{21}&\Sigma_{22}\end{bmatrix}$,
@@
Next, we compute $\Sigma_{21}$ as follows:
$L_{11}^{T}\Sigma_{21}^{T}+L_{21}^{T}\Sigma_{22}=0,\quad\Sigma_{21}=-\Sigma_{22}L_{21}L_{11}^{-1}.$
@@
$\begin{array}[]{rl}\Sigma_{22}&=L_{22}^{-T}L_{22}^{-1},\\[6.0pt] \Sigma_{21}&=-\Sigma_{22}L_{21}L_{11}^{-1},\\[6.0pt] \Sigma_{11}&=-L_{11}^{-T}L_{21}^{T}\Sigma_{21}+L_{11}^{-T}L_{11}^{-1},\\[6.0pt] \Sigma_{20}&=-\Sigma_{21}L_{10}L_{00}^{-1}-\Sigma_{22}L_{20}L_{00}^{-1},\\[6.0pt] \Sigma_{10}&=-\Sigma_{11}L_{10}L_{00}^{-1}-\Sigma_{12}L_{20}L_{00}^{-1},\\[6.0pt] \Sigma_{00}&=-L_{00}^{-T}L_{10}^{T}\Sigma_{10}-L_{00}^{-T}L_{20}^{T}\Sigma_{20}+L_{00}^{-T}L_{00}^{-1}.\end{array}$
@@
1:Initialization:
2:int i, j, k;
3:for i=N1i=N-1 to 0 step 1-1 do
4:  for j=N1j=N-1 to $i$ step 1-1 do
5:   if $i==j$ then
6:     $\Sigma_{ii}\leftarrow\Sigma_{ii}+L_{ii}^{-T}L_{ii}^{-1}$
7:     for $k=i+1$ to NN do
8:      $\Sigma_{ii}\leftarrow\Sigma_{ii}-\Sigma_{ik}L_{ki}L_{ii}^{-1}$
9:     end for
10:   else
11:     $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jj}L_{ji}L_{ii}^{-1}$
12:     for $k=i+1$ to NN do
13:      if $j!=k$ then
14:        $\Sigma_{ji}\leftarrow\Sigma_{ji}-\Sigma_{jk}L_{ki}L_{ii}^{-1}$
15:      end if
16:     end for
17:   end if
18:  end for
19:end for
The core tile operations required for recursive inversion are implemented using well-established matrix computational kernels. These operations are as follows:
@@
  • LAUUM (Lower Triangular Matrix Multiplication): This kernel updates a diagonal tile. It takes the lower triangular part of the tile (denoted as $L_{ii}$) and computes its product with its transpose. The algorithm then mirrors the result to make the tile fully symmetric.
    $\Sigma_{ii}\leftarrow L_{ii}L_{ii}^{T}.$
@@
III-B Selected Elements of the Inverse
Building upon the tile-based sparse Cholesky factorization implemented in sTiles [13], our selected inversion algorithm adapts this concept by pruning all computations not required for the user-specified inverse elements. The overall process is structured into the following key steps:
@@ III-B Selected Elements of the Inverse
    The selection step begins with the user specifying a list of matrix elements, identified by their indices $(i,j)$, for which the inverse is required. These indices are mapped to their corresponding tiles using the same compressed tile format employed during the Cholesky decomposition.
@@
Figure 2: Illustration of Cholesky factorization and inversion patterns of selected tiles for dense (1-5) and arrowhead (6-10) symmetric matrices.
@@
In cases 1-3, the original matrix is fully dense, and the selected tiles include the diagonal. As a result, the inversion covers the entire lower triangle, producing a full inverse. This behavior reflects the high computational cost of inverting dense matrices when the diagonal is involved. Case 6 mirrors this behavior in the arrowhead setting. Although the original structure is sparse, selecting the entire matrix leads to full inversion, analogous to case 1.
@@
III-C Directed Acyclic Graph
@@
The Directed Acyclic Graph (DAG) captures dependencies between computations, enabling efficient parallel execution and optimized resource allocation. By organizing tasks with well-defined precedence, the DAG ensures that independent computations can proceed concurrently. The DAG underlying our approach is constructed using four key computational kernels—TRSM, LAUUM, GEMM, and TRMM—each applied to specific tiles $\Sigma_{i,j}$, corresponding to the tile at position $(i,j)$. The definitions and roles of these operations are described in detail in Section III-A.
@@
As illustrated in Figures 3 and 4, the DAGs for full and arrowhead matrix inversions (Cases 2 and 7, respectively) exhibit notable differences in structure and parallelism. The DAG for full matrix inversion is characterized by a large width, indicating many tasks that could, in principle, be executed concurrently. In the case of the arrowhead structure, the number of nodes is significantly reduced. This visual difference is a direct result of the filtering mechanism mentioned in Section III-A, which prunes unnecessary computations by evaluating only the dependencies required for the user-selected tiles. This leads to a more compact DAG with a reduced overall computational workload. Crucially, the critical path length is unaffected, remaining at six sequential operations along the longest dependency chain for the studied matrices in both the full and arrowhead cases.
Figure 3: Directed acyclic graph (DAG) for selected inversion on a dense matrix of size $6\times 6$ tiles, corresponding to Case 2 in Figure 2. The DAG’s width illustrates the high degree of parallelism available. Its height represents the length of the critical path, which dictates the minimum execution time.
Figure 4: Directed acyclic graph (DAG) for selected inversion on an arrowhead matrix of size $6\times 6$ tiles, corresponding to Case 7 in Figure 2. Compared to the dense case in Figure 3, this DAG has significantly fewer nodes, demonstrating the computational savings from exploiting the matrix structure, while the critical path length remains the same.
III-D Two-Phase Algorithm for Selected Inversion
@@
To parallelize Algorithm 1, we adopt a two-phase approach designed to minimize interdependencies between cores, thereby reducing idle times and enhancing parallel efficiency. This approach serves as the practical implementation of the filtering mechanism, as it prunes unnecessary computations by structuring its loops to operate only on tiles that contribute to the user-specified subset of the inverse. By leveraging the structure of the selected tiles, we avoid redundant operations while preserving correctness. The division of the algorithm into distinct phases enables better task organization and facilitates scalable parallel execution. A critical aspect of this strategy is the use of static load balancing, where tasks are preassigned to cores during preprocessing. This ensures an even distribution of the workload and significantly reduces runtime scheduling overhead.
Phase 1: Independent Row-wise Updates of $L^{T}$
In this phase, we begin the inversion of the upper triangular factor $L^{T}$ by performing a set of computationally independent row-wise operations. The parallelization is achieved by distributing the tile rows of $L^{T}$ among the available cores. The tile rows are assigned in a static, round-robin manner: a core with a given thread ID is responsible for processing the set of block rows indexed by $i$, starting from $i=N-1-\texttt{thread ID}$ and decrementing by total_cores in each step.
This distribution scheme is highly efficient as the computations for different tile rows are completely independent of one another. For any given tile row $i$, all operations only require data from tiles within that same row (i.e., $L_{ii}^{T}$ and $L_{ij}^{T}$). Therefore, cores can proceed concurrently without any need for communication or synchronization during this phase.
The work for each assigned row $i$ consists of two steps, as detailed in Algorithm 2:
  1. 1.
    Diagonal Tile Inversion: The diagonal tile $L_{ii}^{T}$ is inverted. The result, $(L_{ii}^{T})^{-1}$, is stored in the corresponding diagonal block of the output matrix, $\Sigma_{ii}^{T}$. This corresponds to the TRSM operation in the algorithm.
  2. 2.
    Off-Diagonal Tile Update: The newly computed $(L_{ii}^{T})^{-1}$ is then used to update all non-zero off-diagonal tiles $L_{ij}^{T}$ (for $j>i$) in the same block row. This update, performed by the TRMM operation, modifies the $L^{T}$ matrix in-place, preparing it for the second phase of the algorithm.
@@
To handle sparse matrices, we use the notation $j\in\text{neighbors}(i)$ to indicate that an operation is only performed if the tile is structurally non-zero. A non-zero tile may have an initial value of zero but is filled during the symbolic factorization phase and is fully updated by the end of the computation. The algorithm for this phase is detailed in Algorithm 2.
@@
1:Given: Matrix AA is partitioned into $N\times N$ tiles and expressed as $A=LL^{T}$, where $L^{T}$ is the upper triangular factor.
2:Note: $\Sigma$ is the matrix where the selected inverse is stored.
3:Initialization: $i\leftarrow N-1-\texttt{thread ID}$
4:while $i\geq 0$ do
5:  for all $j\in\text{neighbors}(i)$ and $i\leq j<N-1$ do
6:   if $i==j$ then
7:     $\Sigma_{ii}^{T}\leftarrow\texttt{TRSM}(L_{ii}^{T},I)$
8:   else
9:     $L_{ij}^{T}\leftarrow\texttt{TRMM}(L_{ii}^{T},L_{ij}^{T})$
10:   end if
11:  end for-
12:  $i\leftarrow i-\texttt{total\_cores}$
13:end while
@@
Once Phase 1 is complete, the algorithm transitions to a second phase to finalize the selected inverse, $\Sigma$. Unlike the first phase, this stage involves complex data dependencies between tile computations. The calculation of a tile $\Sigma_{ij}$ often requires results from other tiles (e.g., $\Sigma_{kj}$), which may be computed by different cores. These dependencies create a task graph where cores must synchronize to ensure correct computational ordering. To manage these dependencies, we implement a lightweight, asynchronous producer-consumer model. This model is realized using a shared status-tracking matrix, core_progress, which allows threads to signal the completion of a tile and wait for dependencies without expensive global barriers. The producer-consumer mechanism works as follows:
  • Producer Role: When a core successfully computes a tile $\Sigma_{xy}$, it acts as a producer by updating the corresponding entry in core_progress to a ”completed” state. This update signals to all other cores that the data in $\Sigma_{xy}$ is now ready to be used.
  • Consumer Role: Before a core computes a tile that depends on $\Sigma_{xy}$, it acts as a consumer. It polls the status of the prerequisite tile by reading the core_progress matrix, waiting until it is marked as complete.
For clarity in the pseudocode (Algorithm 3), we abstract this mechanism into two high-level primitives:
  • WaitForTile(row, col): Represents the consumer’s action of waiting for the flag in core-progress to be set.
  • SignalTileReady(row, col): Represents the producer’s action of setting the flag upon task completion.
This approach allows the algorithm to focus on the logical dataflow, enabling a high degree of parallelism as cores can work on any available tasks whose dependencies have been met.
@@
1:Define synchronization primitives:
2:WaitForTile(row, col): Pauses until tile $\Sigma_{\texttt{row, col}}$ is marked complete.
3:SignalTileReady(row, col): Marks tile $\Sigma_{\texttt{row, col}}$ as complete.
4:
5:Initialization: $i\leftarrow N-1$- thread ID and set = false;
6:while $i\geq 0$ do
7:  for j=N1j=N-1 to $i$ step 1-1 do
8:   if $i==j$ then
9:     $\Sigma_{ii}\leftarrow\texttt{LAUUM}(\Sigma_{ii})$
10:     for all $k\in\text{neighbors}(i)$ and i<k<N1i<k<N-1 do
11:      WaitForTile(i, k)
12:      $\Sigma_{ii}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{ik}^{T},\Sigma_{ii})$
13:     end for
14:     SignalTileReady(i,i)
15:   else
16:     if $(i+1)\leq(N-1)$ then
17:      set = true; $i$ = 0; $j$ = 0;
18:     end if
19:     for all $k\in\text{neighbors}(j)$ and i<k<N1i<k<N-1 do
20:      if $i\in\text{neighbors}(j)\cap j\in\text{neighbors}(k)$ and $i<j$ then
21:        if $k>j$ then
22:         WaitForTile(j, k)
23:         $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{kj},\Sigma_{ij})$
24:         $i$ = i; $j$ = j;
25:        else
26:          WaitForTile(k, j)
27:         $\Sigma_{ij}\leftarrow\texttt{GEMM}(L_{ik},\Sigma_{jk}^{T},\Sigma_{ij})$
28:         $i$ = i; $j$ = j;
29:        end if
30:      end if
31:     end for
32:     if set then
33:      SignalTileReady(ii, $j$)
34:     end if
35:   end if
36:  end for
37:  $i\leftarrow i-\texttt{total\_cores}$
38:end while